OpenStudy (anonymous):

You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this information? y = 0.03x + 15 y = 0.03x - 15 y = -0.03x + 15 y = -0.03x - 15

OpenStudy (anonymous):

Let y = gallons of gasoline Let x = number of miles y starts at 15 and then you subtract 0.03 gal. for each x.

OpenStudy (anonymous):

so its the first one?

OpenStudy (anonymous):

Does the first one include the subtraction I mentioned?

OpenStudy (anonymous):

y=15-.03x

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@akash_809 http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

when x=0 then y=15 and with every minute y decreases by 0.03 , so u get y=15-0.03x

OpenStudy (anonymous):

@CliffSedge u a mod

OpenStudy (anonymous):

no but im using y=mx+b which means m=what could change which is .03 and then b=15 because it is your priginal so shouldnt it be y=.03x+15

OpenStudy (anonymous):

@akash_809 does that matter?

OpenStudy (anonymous):

@EmmaH do u get it

OpenStudy (anonymous):

m is negative as y is decreasing

OpenStudy (anonymous):

it cant b that because it is increasing because it increases as the cars goes

OpenStudy (anonymous):

could you try to explain step by step please

OpenStudy (anonymous):

what is slope of a line or any function y=f(x)...it is m=dy/dx i.e rate of change of y wrt x and here rate of change of y i.e fuel is negative w.r.t x...now is it better

OpenStudy (anonymous):

i know i must sound stupid but that doesnt make sense to me sorry